A function is defined by the polynomial 𝑓(π‘₯) = 3π‘₯ 4 βˆ’ 4π‘₯ 3 βˆ’ 12π‘₯ 2 + 8. Find and classify all the stationary points of f(x).


1
Expert's answer
2021-06-08T17:21:18-0400

"f(x)=3x^4-4x^3-12x^2+8"

differentiating with respect to x

"f'(x)=12x^3-12x^2-24x"

to find stationary points put "f'(x)=0"

"\Rightarrow12x^3-12x^2-24x=0"

"\Rightarrow x^3-x^2-2x=0"

"\Rightarrow x(x^2-x-2)=0"

"\Rightarrow x(x-2)(x+1)=0"


stationary points are

"x=-1,0,2"


differentiating again with respect to x

"f''(x)=36x^2-24x-24"

"f''(-1)=36>0 \space(x=-1 \space is \space local\space minimum)"

"f''(0)=-24<0\space(x=0 \space is \space local \space maximum)"

"f''(2)=72>0\space(x=2 \space is \space local\space minimum)"